OpenStudy (anonymous):

if a+b=s and a-b=t, then which of the following expresses the value of ab in terms of s and t? a. st b. (s-t)/2 c. (s+t)/2 d. (s^2-t^2)/4 e. (s^2-t^2)/2

OpenStudy (littlebird):

You just have to check each one. For example: s-t/2 = ((a+b)-(a-b))/2 = 2b/2= b This means that answer b is wrong, and for similar reasons, c is also wrong. You can tell a is wrong just by looking at it. d and e are very similar, so I only have to do one more test, which is in the attached image.